Mia's motorboat took 3 hours to make a trip downstream with a 6 mph current. the return trip against the came current took 5 hours. find the speed of the boat.

please help me thank you!

Answer by stanbon(75887)

Mia's motorboat took 3 hours to make a trip downstream with a 6 mph current. the return trip against the came current took 5 hours. find the speed of the boat.
-----------------
down Data:
time = 3 hrs ; rate = (b+6) mph ; distance = rt = 3(r+6) miles
--------------------
upstream DATA:
time = 5 hrs ; rate = (b-6) mph ; distance = rt = 5(r-6) miles
--------------------------------------------------------------
Equation:
distance up = distance down
5(b-6) = 3(b+6)
5b - 30 = 3b + 18
2b = 48
b = 24 mph (speed of the boat in still water)
==================
Cheers,
Stan H.